OpenStudy (anonymous):

Will fan and medal :) Enter the answer. 5. What value of k makes the equation true? 12 + (–17) = –17 + k k = ____

OpenStudy (cypress7):

the community property of addition stated that if A+B then B+A so whats the answer?

OpenStudy (yamyam70):

Solve fir K first

OpenStudy (yamyam70):

then substitute k nto the original question

OpenStudy (yamyam70):

if both sides are equal, then the value of K is true.

OpenStudy (dtan5457):

just looking at 12 + (–17) = –17 + k i think it's evident k=12 it has to be the same as the other side

OpenStudy (yamyam70):

it should be same both sides, so what did yoiu get?

OpenStudy (yamyam70):

so pplug 12 as K

OpenStudy (yamyam70):

12+ (-17) = -17 +12 ????

OpenStudy (dtan5457):

-5=-5?

OpenStudy (anonymous):

ya the it was 12 thank you guys :)